OpenStudy (anonymous):

Here is another... A tv screen is 2 inches longer than it's height. If the diagonal length of the screen is 2 inches more than it's length, what is the diagonal measurement of the screen?

OpenStudy (anonymous):

put height = x, length = x+2 and diagonal = x+4 then use pythatoras \[x^2+(x+2)^2=(x+4)^2\]

OpenStudy (anonymous):

\[x^2+x^2+4x+4=x^2+8x+16\] \[2x^2+4x+4=x^2+8x+16\] \[x^2-4x-12=0\]

OpenStudy (anonymous):

\[(x-6)(x+2)=0\] \[x-6=0\] \[x=6\] we can ignore the other answer because it is negative

OpenStudy (anonymous):

awesome thank you!!!